### v2

@@ 4.2 Deriving the strong second-order sufficient condition
Here we reformulate the second-order optimality conditions (4.5) and (4.6) to fit (4.17) in Corollary 4.1.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (4.12), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then the second-order sufficient condition (4.5) is equivalent to
$\langle{\bm{\nu}},H^{\top}\Theta H{\bm{\nu}}\rangle>0,\quad\forall\,{\bm{\nu}}\in{\mathcal{V}}\setminus\{{\boldsymbol{0}}\},$ (4.21)
@@
${\mathcal{V}}=\left\{{\bm{\nu}}\in{\mathds{R}}^{\ell}\,\middle|\,{\mathcal{G}}_{j}{\bm{\nu}}\geq 0,\,\forall\,j\in J_{4,5},\ {\mathcal{G}}_{j}{\bm{\nu}}\in{\mathcal{Q}}_{j},\,\forall\,j\in J_{6}\right\}$ (4.22)
with each ${\mathcal{G}}_{j}$ being defined in (4.18). Moreover, the strong second-order sufficient condition (4.6) is equivalent to $H^{\top}\Theta H\succ O$.
Proof.
According to the definition of $\Theta$ in (4.12), one can rewrite (4.5) as
$\langle{\boldsymbol{d}},\Theta{\boldsymbol{d}}\rangle>0,\quad\forall\,{\boldsymbol{d}}\in{\mathcal{C}}({\boldsymbol{x}}^{*})\setminus\{{\boldsymbol{0}}\},$ (4.23)
where the critical cone ${\mathcal{C}}({\boldsymbol{x}}^{*})$ is given by (4.3). Since $H\in{\mathds{R}}^{n\times\ell}$ is of full column rank and ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$, from the definition of the critical cone in (4.3) and its affine hull in (4.4), one can see that the second-order sufficient condition (4.23) is equivalent to (4.21) with
@@
Based on the explicit formulations of the tangent cones in (4.2), one can reformulate ${\mathcal{V}}$ as the one in (4.22). The conclusion on the strong second-order sufficient condition directly comes from the fact that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. This completes the proof. ∎
Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (4.12). Suppose that Assumption 4.1 holds and that $H\in{\mathds{R}}^{n\times\ell}$ is a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. To prove the strong second-order sufficient condition, i.e., $H^{\top}\Theta H\succ O$, we construct a series of symmetric matrices of different dimensions and prove that they are positive definite.
We further suppose that (4.17) holds with ${\mathcal{G}}_{j}$ ($j\in J_{4,5,6}$) being defined by (4.18). For convenience, for each integer $i\in\{1,\ldots,|J_{4,5,6}|\}$, we assign it a unique index $j\in J_{4,5,6}$ denoted by $j_{i}$. Without loss of generality, we assume that
$j_{1},\ldots,j_{|J_{4}|}\subseteq J_{4}\quad\mbox{and}\quad j_{|J_{4}|+1},\ldots,j_{|J_{4,5}|}\subseteq J_{5}.$ (4.24)
Then one can write (4.18) as follows
${\mathcal{G}}_{j_{i}}:=\begin{cases}((\dot{\mu}^{*})^{j_{i}};-(\bar{\bm{\mu}}^{*})^{j_{i}})^{\top}{\mathcal{J}}g^{j_{i}}({\boldsymbol{x}}^{*})H,&i=1,\ldots,|J_{4}|,\\ (\dot{g}^{j_{i}}({\boldsymbol{x}}^{*});-\bar{g}^{j_{i}}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j_{i}}({\boldsymbol{x}}^{*})H,&i=|J_{4}|+1,\ldots,|J_{4,5}|,\\ {\mathcal{J}}g^{j_{i}}({\boldsymbol{x}}^{*})H,&i=|J_{4,5}|+1,\ldots,|J_{4,5,6}|.\end{cases}$ (4.25)
Recall from Corollary 4.1 that $H^{\top}\Theta H$ is nonsingular and (4.20) holds. Based on (4.25), we can define the following matrices recursively:
$\begin{cases}R_{i}\in{\mathds{R}}^{(\ell-i+1)\times(\ell-i)},&i=1,\ldots,|J_{4,5}|,\\ R_{i}\in{\mathds{R}}^{\left(\ell-i+1-\sum_{k=|J_{4,5}|+1}^{i-1}r_{j_{k}}\right)\times\left(\ell-i-\sum_{k=|J_{4,5}|+1}^{i}r_{j_{k}}\right)},&i=|J_{4,5}|+1,\,\ldots\,,|J_{4,5,6}|\end{cases}$ (4.26)
@@
$\begin{cases}P_{i}:=R_{1}\times R_{2}\times\cdots\times R_{i}\in{\mathds{R}}^{\ell\times(\ell-i)},&i=1,\ldots,|J_{4,5}|,\\ P_{i}:=R_{1}\times R_{2}\times\cdots\times R_{i}\in{\mathds{R}}^{\ell\times\left(\ell-i-\sum_{k=|J_{4,5}|+1}^{i}r_{j_{k}}\right)},&i=|J_{4,5}|+1,\ldots,|J_{4,5,6}|.\end{cases}$ (4.28)
@@
$T_{i}:=P_{i}^{\top}H^{\top}\Theta HP_{i}=R_{i}^{\top}\times\cdots\times R_{1}^{\top}T_{0}R_{1}\times\cdots\times R_{i},\quad i=1,\ldots,|J_{4,5,6}|.$ (4.29)
It is obvious from (4.27) that each $P_{i}$ has full column rank. Moreover, we have the following result, which is the most crucial step in our analysis.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (4.12) and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Suppose that (4.17) holds. For the matrices defined in (4.26), (4.28) and (4.29), one has that each $T_{i}$ is nonsingular, and
$[{\mathcal{G}}_{j_{i}}P_{i-1}]T_{i-1}^{-1}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top}\succ O,\quad\forall\,i=1,\ldots,|J_{4,5,6}|.$ (4.30)
Proof.
Since (4.17) holds in this setting, by using (4.24) we can get that, for any ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$,
$\begin{array}[]{ll}H^{\top}\Theta H{\bm{\nu}}\notin-\left\{\sum\limits_{i=1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}p^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \Big{|}\ \begin{array}[]{ll}p^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}{\bm{\nu}}),\\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}{\bm{\nu}})\end{array}\right\}.\end{array}$ (4.31)
@@
$\begin{array}[]{ll}T_{k}{\bm{\eta}}\notin-P_{k}^{\top}\left\{\sum\limits_{i=k+1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}p^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\Big{|}\begin{array}[]{ll}p^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}),\\[2.84526pt] {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\end{array}\right\}.\end{array}$ (4.32)
Recall that $T_{0}=H^{\top}\Theta H$. Then, by (4.31) we know that (4.32) holds for $k=0$. If $1\leq k\leq|J_{4,5}|$, one has from (4.28) that ${\mathcal{G}}_{j_{i}}P_{k}={\mathcal{G}}_{j_{i}}R_{1}\times R_{2}\times\cdots\times R_{k}$. Then one can get from (4.27) that ${\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}=0$ for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-k}$ and all $i=1,\ldots,k$. Thus, by (2.4) we know that (4.31) implies that, for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-k}\setminus\{{\boldsymbol{0}}\}$, it holds that
$\begin{array}[]{ll}H^{\top}\Theta HP_{k}{\bm{\eta}}\notin{\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right)\\[5.69054pt] \qquad-\left\{\sum\limits_{i=k+1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}p^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \Big{|}\ \begin{array}[]{ll}p^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}),\\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\end{array}\right\}.\end{array}$ (4.33)
Note that $P_{k}^{\top}{\mathcal{G}}_{j_{i}}^{\top}={\boldsymbol{0}}$ for all $i=1,\ldots,k$, so that $\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\subseteq\ker P_{k}^{\top}$. Moreover, as $P_{k}^{\top}\in{\mathds{R}}^{(l-k)\times l}$ has full row rank by definition, the dimension of $\ker P_{k}^{\top}$ is exactly $k$. Then by Corollary 4.2 we know that ${\sf span}(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k})=\ker P_{k}^{\top}$. Suppose to the contrary that (4.32) can not be true. That is there exist a nonzero vector $\hat{\bm{\eta}}\in{\mathds{R}}^{\ell-k}$, scalars $\hat{p}^{j_{i}}\in{\mathds{R}}$ for all $i=k+1,\ldots,|J_{4,5}|$, and vectors $\hat{\boldsymbol{p}}^{j_{i}}\in{\mathds{R}}^{1+r_{j_{i}}}$ for all $i=|J_{4,5}|+1,\ldots,|J_{4,5,6}|$ satisfying $\hat{p}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}P_{k}\hat{\bm{\eta}})$ and $\hat{\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}\hat{\bm{\eta}})$, such that
@@
$\begin{array}[]{ll}H^{\top}\Theta HP_{k}\hat{\bm{\eta}}+\sum\limits_{i=k+1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}\hat{p}^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}\hat{\boldsymbol{p}}^{j_{i}}\in\ker P_{k}^{\top}={\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right),\end{array}$
which contradicts (4.33). Particularly, if $T_{k}\hat{\bm{\eta}}={\boldsymbol{0}}$, one can take $\hat{p}^{j_{i}}=0$ and $\hat{\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ due to (2.4) and Part (6e) of Lemma 3.2. Therefore, we know that (4.32) holds and $T_{k}$ is nonsingular.
@@
$\begin{array}[]{ll}T_{k}{\bm{\eta}}\notin-P_{k}^{\top}\left\{\sum\limits_{i=k+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \big{|}\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\right\},\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}\setminus\{{\boldsymbol{0}}\},\end{array}$ (4.34)
where $\tilde{k}:=k+\sum_{t=|J_{4,5}|+1}^{k}r_{j_{t}}$. If $|J_{4,5}|+1\leq k\leq|J_{4,5,6}|$, from (4.28) one has ${\mathcal{G}}_{j_{i}}P_{k}={\mathcal{G}}_{j_{i}}R_{1}\times R_{2}\times\cdots\times R_{k}$. According to (4.27) one can get ${\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}={\boldsymbol{0}}$ for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}$ and any $i=1,\ldots,k$. Then, by (2.4) and Part (6a) of Lemma 3.2 (or (2.4) if $r_{j}=0$) we know that (4.31) implies that, for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}\setminus\{{\boldsymbol{0}}\}$,
$\begin{array}[]{ll}H^{\top}\Theta HP_{k}{\bm{\eta}}\notin{\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right)-\left\{\sum\limits_{i=k+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \Big{|}\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\right\}.\end{array}$ (4.35)
@@
$\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,|J_{4,5}|}\cup\{{\sf span}{\mathcal{G}}_{j_{i}}^{\top}\}_{i=|J_{4,5}|+1,\ldots,k}\subseteq\ker P_{k}^{\top}.$
Then by Corollary 4.2 we know that ${\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right)=\ker P_{k}^{\top}$. Suppose to the contrary that (4.34) is not true, i.e., there exists a nonzero vector $\tilde{\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}$ and vectors $\tilde{\boldsymbol{p}}^{j_{i}}\in{\mathds{R}}^{1+r_{j_{i}}}$ ($i=k+1,\ldots,|J_{4,5,6}|$) such that

### v3

@@ 4.2 Deriving the strong second-order sufficient condition
Here we reformulate the second-order optimality conditions (31) and (32) to fit (43) in Corollary 4.9.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (38), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then the second-order sufficient condition (31) is equivalent to
(47) $\langle{\bm{\nu}},H^{\top}\Theta H{\bm{\nu}}\rangle>0,\quad\forall\,{\bm{\nu}}\in{\mathcal{V}}\setminus\{{\boldsymbol{0}}\},$
@@
(48) ${\mathcal{V}}=\left\{{\bm{\nu}}\in{\mathds{R}}^{\ell}\,\middle|\,{\mathcal{G}}_{j}{\bm{\nu}}\geq 0,\,\forall\,j\in J_{4,5},\ {\mathcal{G}}_{j}{\bm{\nu}}\in{\mathcal{Q}}_{j},\,\forall\,j\in J_{6}\right\}$
with each ${\mathcal{G}}_{j}$ being defined in (44). Moreover, the strong second-order sufficient condition (32) is equivalent to $H^{\top}\Theta H\succ O$.
Proof 4.14.
According to the definition of $\Theta$ in (38), one can rewrite (31) as
(49) $\langle{\boldsymbol{d}},\Theta{\boldsymbol{d}}\rangle>0,\quad\forall\,{\boldsymbol{d}}\in{\mathcal{C}}({\boldsymbol{x}}^{*})\setminus\{{\boldsymbol{0}}\},$
where the critical cone ${\mathcal{C}}({\boldsymbol{x}}^{*})$ is given by (29). Since $H\in{\mathds{R}}^{n\times\ell}$ is of full column rank and ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$, from the definition of the critical cone in (29) and its affine hull in (30), one can see that the second-order sufficient condition (49) is equivalent to (47) with
@@
Based on the explicit formulations of the tangent cones in (28), one can reformulate ${\mathcal{V}}$ as the one in (48). The conclusion on the strong second-order sufficient condition directly comes from the fact that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. This completes the proof.
Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (38). Suppose that Assumption 4.1 holds and that $H\in{\mathds{R}}^{n\times\ell}$ is a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. To prove the strong second-order sufficient condition, i.e., $H^{\top}\Theta H\succ O$, we construct a series of symmetric matrices of different dimensions and prove that they are positive definite.
We further suppose that (43) holds with ${\mathcal{G}}_{j}$ ($j\in J_{4,5,6}$) being defined by (44). For convenience, for each integer $i\in\{1,\ldots,|J_{4,5,6}|\}$, we assign it a unique index $j\in J_{4,5,6}$ denoted by $j_{i}$. Without loss of generality, we assume that
(50) $j_{1},\ldots,j_{|J_{4}|}\subseteq J_{4}\quad\mbox{and}\quad j_{|J_{4}|+1},\ldots,j_{|J_{4,5}|}\subseteq J_{5}.$
Then one can write (44) as follows
(51) ${\mathcal{G}}_{j_{i}}:=\begin{cases}((\dot{\mu}^{*})^{j_{i}};-(\bar{\bm{\mu}}^{*})^{j_{i}})^{\top}{\mathcal{J}}g^{j_{i}}({\boldsymbol{x}}^{*})H,&i=1,\ldots,|J_{4}|,\\ (\dot{g}^{j_{i}}({\boldsymbol{x}}^{*});-\bar{g}^{j_{i}}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j_{i}}({\boldsymbol{x}}^{*})H,&i=|J_{4}|+1,\ldots,|J_{4,5}|,\\ {\mathcal{J}}g^{j_{i}}({\boldsymbol{x}}^{*})H,&i=|J_{4,5}|+1,\ldots,|J_{4,5,6}|.\end{cases}$
Recall from Corollary 4.9 that $H^{\top}\Theta H$ is nonsingular and (46) holds. Based on (51), we can define the following matrices recursively:
(52) $\begin{cases}R_{i}\in{\mathds{R}}^{(\ell-i+1)\times(\ell-i)},&i=1,\ldots,|J_{4,5}|,\\ R_{i}\in{\mathds{R}}^{\left(\ell-i+1-\sum_{k=|J_{4,5}|+1}^{i-1}r_{j_{k}}\right)\times\left(\ell-i-\sum_{k=|J_{4,5}|+1}^{i}r_{j_{k}}\right)},&i=|J_{4,5}|+1,\,\ldots\,,|J_{4,5,6}|\end{cases}$
@@
(54) $\begin{cases}P_{i}:=R_{1}\times R_{2}\times\cdots\times R_{i}\in{\mathds{R}}^{\ell\times(\ell-i)},&i=1,\ldots,|J_{4,5}|,\\ P_{i}:=R_{1}\times R_{2}\times\cdots\times R_{i}\in{\mathds{R}}^{\ell\times\left(\ell-i-\sum_{k=|J_{4,5}|+1}^{i}r_{j_{k}}\right)},&i=|J_{4,5}|+1,\ldots,|J_{4,5,6}|.\end{cases}$
@@
(55) $T_{i}:=P_{i}^{\top}H^{\top}\Theta HP_{i}=R_{i}^{\top}\times\cdots\times R_{1}^{\top}T_{0}R_{1}\times\cdots\times R_{i},\quad i=1,\ldots,|J_{4,5,6}|.$
It is obvious from (53) that each $P_{i}$ has full column rank. Moreover, we have the following result, which is the most crucial step in our analysis.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (38) and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Suppose that (43) holds. For the matrices defined in (52), (54) and (55), one has that each $T_{i}$ is nonsingular, and
(56) $[{\mathcal{G}}_{j_{i}}P_{i-1}]T_{i-1}^{-1}[{\mathcal{G}}_{j_{i}}P_{i-1}]^{\top}\succ O,\quad\forall\,i=1,\ldots,|J_{4,5,6}|.$
Proof 4.16.
Since (43) holds in this setting, by using (50) we can get that, for any ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$,
(57) $\begin{array}[]{ll}H^{\top}\Theta H{\bm{\nu}}\notin-\left\{\sum\limits_{i=1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}p^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \Big{|}\ \begin{array}[]{ll}p^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}{\bm{\nu}}),\\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}{\bm{\nu}})\end{array}\right\}.\end{array}$
@@
(58) $\begin{array}[]{ll}T_{k}{\bm{\eta}}\notin-P_{k}^{\top}\left\{\sum\limits_{i=k+1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}p^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\Big{|}\begin{array}[]{ll}p^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}),\\[2.84526pt] {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\end{array}\right\}.\end{array}$
Recall that $T_{0}=H^{\top}\Theta H$. Then, by (57) we know that (58) holds for $k=0$. If $1\leq k\leq|J_{4,5}|$, one has from (54) that ${\mathcal{G}}_{j_{i}}P_{k}={\mathcal{G}}_{j_{i}}R_{1}\times R_{2}\times\cdots\times R_{k}$. Then one can get from (53) that ${\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}=0$ for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-k}$ and all $i=1,\ldots,k$. Thus, by (12) we know that (57) implies that, for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-k}\setminus\{{\boldsymbol{0}}\}$, it holds that
(59) $\begin{array}[]{ll}H^{\top}\Theta HP_{k}{\bm{\eta}}\notin{\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right)\\[5.69054pt] \qquad-\left\{\sum\limits_{i=k+1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}p^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \Big{|}\ \begin{array}[]{ll}p^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}),\\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\end{array}\right\}.\end{array}$
Note that $P_{k}^{\top}{\mathcal{G}}_{j_{i}}^{\top}={\boldsymbol{0}}$ for all $i=1,\ldots,k$, so that $\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\subseteq{\sf ker}P_{k}^{\top}$. Moreover, as $P_{k}^{\top}\in{\mathds{R}}^{(l-k)\times l}$ has full row rank by definition, the dimension of ${\sf ker}P_{k}^{\top}$ is exactly $k$. Then by Corollary 4.11 we know that ${\sf span}(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k})={\sf ker}P_{k}^{\top}$. Suppose to the contrary that (58) can not be true. That is there exist a nonzero vector $\hat{\bm{\eta}}\in{\mathds{R}}^{\ell-k}$, scalars $\hat{p}^{j_{i}}\in{\mathds{R}}$ for all $i=k+1,\ldots,|J_{4,5}|$, and vectors $\hat{\boldsymbol{p}}^{j_{i}}\in{\mathds{R}}^{1+r_{j_{i}}}$ for all $i=|J_{4,5}|+1,\ldots,|J_{4,5,6}|$ satisfying $\hat{p}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j_{i}}P_{k}\hat{\bm{\eta}})$ and $\hat{\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}\hat{\bm{\eta}})$, such that
@@
$\begin{array}[]{ll}H^{\top}\Theta HP_{k}\hat{\bm{\eta}}+\sum\limits_{i=k+1}^{|J_{4,5}|}{\mathcal{G}}_{j_{i}}^{\top}\hat{p}^{j_{i}}+\sum\limits_{i=|J_{4,5}|+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}\hat{\boldsymbol{p}}^{j_{i}}\in{\sf ker}P_{k}^{\top}={\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right),\end{array}$
which contradicts (59). Particularly, if $T_{k}\hat{\bm{\eta}}={\boldsymbol{0}}$, one can take $\hat{p}^{j_{i}}=0$ and $\hat{\boldsymbol{p}}^{j_{i}}={\boldsymbol{0}}$ due to (12) and Part (6e) of Lemma 3.2. Therefore, we know that (58) holds and $T_{k}$ is nonsingular.
@@
(60) $\begin{array}[]{ll}T_{k}{\bm{\eta}}\notin-P_{k}^{\top}\left\{\sum\limits_{i=k+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \big{|}\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\right\},\forall\,{\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}\setminus\{{\boldsymbol{0}}\},\end{array}$
where $\tilde{k}:=k+\sum_{t=|J_{4,5}|+1}^{k}r_{j_{t}}$. If $|J_{4,5}|+1\leq k\leq|J_{4,5,6}|$, from (54) one has ${\mathcal{G}}_{j_{i}}P_{k}={\mathcal{G}}_{j_{i}}R_{1}\times R_{2}\times\cdots\times R_{k}$. According to (53) one can get ${\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}}={\boldsymbol{0}}$ for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}$ and any $i=1,\ldots,k$. Then, by (12) and Part (6a) of Lemma 3.2 (or (12) if $r_{j}=0$) we know that (57) implies that, for any ${\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}\setminus\{{\boldsymbol{0}}\}$,
(61) $\begin{array}[]{ll}H^{\top}\Theta HP_{k}{\bm{\eta}}\notin{\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right)-\left\{\sum\limits_{i=k+1}^{|J_{4,5,6}|}{\mathcal{G}}_{j_{i}}^{\top}{\boldsymbol{p}}^{j_{i}}\ \Big{|}\ {\boldsymbol{p}}^{j_{i}}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j_{i}}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j_{i}}P_{k}{\bm{\eta}})\right\}.\end{array}$
@@
$\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,|J_{4,5}|}\cup\{{\sf span}{\mathcal{G}}_{j_{i}}^{\top}\}_{i=|J_{4,5}|+1,\ldots,k}\subseteq{\sf ker}P_{k}^{\top}.$
Then by Corollary 4.11 we know that ${\sf span}\left(\{{\mathcal{G}}_{j_{i}}^{\top}\}_{i=1,\ldots,k}\right)={\sf ker}P_{k}^{\top}$. Suppose to the contrary that (60) is not true, i.e., there exists a nonzero vector $\tilde{\bm{\eta}}\in{\mathds{R}}^{\ell-\tilde{k}}$ and vectors $\tilde{\boldsymbol{p}}^{j_{i}}\in{\mathds{R}}^{1+r_{j_{i}}}$ ($i=k+1,\ldots,|J_{4,5,6}|$) such that